Home‎ > ‎Crew 131 - Embry Riddle‎ > ‎

1216-CommandersReport

Commander Report
12/16/2013
Chelsea Iwig
 
            This morning Crew 131 were able to enjoy their sleep a little longer as they did not need to be up until 8am to prepare for the arrival of a media crew who had traveled from Earth just to interview the current inhabitants of the Mars Hab. After finishing up some last minute chores around the Hab, Crew 131 sat down together to play some card games as they anxiously awaited the arrival of their first guests from Earth since arriving on the Red Planet.
            Finally at around 1030, their waiting was over when they heard knocking on the Hab hatch. Members of the crew rushed down the stairs to open the hatch for their guests and then proceeded through brief introductions before beginning a tour of the Hab. Once the introductions and tour was complete, it was time for the EVA team to suit up for the walking EVA that had been planned for the late morning through the early afternoon. This EVA was intentionally made fairly easy and relaxed as they knew the media crew would need to coordinate filming of the EVA as well. Therefore, the team decided to just return to the gorge a little over 2km East of the Hab that they had visited two days prior. However, this time, upon arrival at their destination, they spent much more time exploring further south where a large outcropping over the gorge provided for some breathtaking views of the Martian surface.
            The media crew was clearly impressed with the amazing views as well as the difficulty and exertion involved in completing even just a one and-a-half hour EVA. When the EVA team finally returned to the Hab, they proceeded to show the media crew how they prepare meals by cooking up one of their favorite meals since arriving on Mars – cheesy potato soup. While some crewmembers worked on cooking the soup and baking some homemade chocolate chip cookies, the media crew proceeded to interview three of the other crewmembers. Once lunch and the interviews were completed, the Media crew returned in their space capsule to the nearby station they would be staying at for the next couple of days to regroup and also to stop on their way at the service station in order to interview the expert technician who helps the crew with mechanical issues from time to time.  
             The media crew plans to return again tonight to conduct some further interviews about what life on Mars is like before returning again to their station for the night. After Crew 131 finishes sending in their reports to Mission Support, they will complete their surveys for the evening and then they would like to visit the Observatory to conduct some star observations if Mission Support clears it. If not, the crew will likely watch a movie or spend time together playing card games until it is time for them to go to bed. Since tomorrow will be an early morning and a very long day, the crew will try to be in bed by 2300 tonight.

-------------------------

CommanderCheckIn Report
12/16/2013
Chelsea Iwig
Crew Physical Health:
Two crewmembers have a cough. All other crewmembers are in good 
physical health. 
   
Eva Depart Time:
1108
Eva Return Time:
1240
Narrative Of Field Mission Results:
Team 2 along with the two media visitors went on a short walking EVA to the same location east of the Hab that Team 2 visited on Saturday; however, this time they explored more to the south of their destination as well. They tested out the armbands and were able to come up with ideas to fix the problem of not being able to read the screen in bright sunlight. These solutions are not things that can be completed during our current session, but hopefully, the changes can be made when we get home for future testing. Overall, the media crew seemed enjoyed the EVA.
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
   Today DG came by and switched to the new generators.
The HVAC system has been running smoothly, no turning off. We believe this may be due to the gas regulators. Still waiting to hear from gas company.
Propane is low -- expected to be delivered tomorrow.
Pipes from the static tank have been wrapped with heat tape. The crew engineer will refrain from turning on the pump through the night to see if it can prevent freezing.
Rep Sched:
  Commander Check-In, Commander Report, Engineering Report, HSO Report, EVA Plan, EVA Report, Journalist Report. 
Tomorrow Plan:
   Tomorrow the crew will be waking up bright and early at 0630 for the high intensity workout again. However, this time, the media crew will be coming back to the Hab to finish up some filming of our exercise study as well as crew interviews. Once Crew 131 has finished the exercise study and the interviews for the media crew, they will begin preparing for Team 1’s EVA. Once Team 1 has returned after their EVA, the crew will complete the glove study and prepare their reports for COMMS. Following COMMS the crew will complete the surveys for the Space Team Assessment Tool and relax for the rest of the evening. 
Inventory:
   None
Support Req:
   New filters for the EVA packs and rat traps are needed. Also, fluorescent light bulbs for the downstairs lab are needed as only one is working currently and it has been flickering on and off. We found some more dish detergent, but we are still low. We will probably be able to make it last until the end of our session, but the next crew will definitely need more. 
Miscelleneous:
   Still no luck catching the rogue mice and rats.